# Searching for internal symbols underlying deep learning

### v1

@@ Abstract
Deep learning (DL) enables deep neural networks (DNNs) to automatically learn complex tasks or rules from given examples without instructions or guiding principles. As we do not engineer DNNs’ functions, it is extremely difficult to diagnose their decisions, and multiple lines of studies proposed to explain principles of DNNs/DL operations. Notably, one line of studies suggests that DNNs may learn concepts, the high level features recognizable to humans. Thus, we hypothesized that DNNs develop abstract codes, not necessarily recognizable to humans, which can be used to augment DNNs’ decision-making. To address this hypothesis, we combined foundation segmentation models and unsupervised learning to extract internal codes and identify potential use of abstract codes to make DL’s decision-making more reliable and safer.
@@ 1 Introduction
Deep learning (DL) can automatically wire deep neural networks (DNNs) to perform complex tasks, and DNNs’ learning capability allows them to outperform humans in some tasks [1, 2, 3]. Notably, instead of expert instructions or feature engineering, DL requires only training examples to solve highly complex problems. However, as DNNs’ structures are automatically determined without guiding principles, we do not fully understand their decision-making process. That is, DNNs may be useful black box decision-makers, but without safety measures, deploying them in high-stakes domains (e.g., autonomous cars) comes with unpredictable risks [4, 5]. For the moment, the current generation of DL does not warrant essential safety to be deployed in safety-critical or high-stakes domains. To address this shortcoming, earlier studies sought potential methods that could decode and explain DNNs’ decisions. There are multiple lines of research in explainable AI (e.g., see Refs. [6, 7, 8, 9, 10]), but the most relevant to this study is the one probing correlations between hidden layer representations and inputs or between hidden layer representations and outputs (i.e., decisions). Feature visualization [11] and network dissection [12] sought direct links between inputs and hidden layers, while linear probes investigated correlations between outputs and hidden layers [13, 14].
@@
It is notable that the mapping between hidden layers and visual features in input layers, in principle, establishes links between high level patterns recognizable to humans (often referred to as ‘concepts’) and hidden neurons. The Test with Concept Activation Vectors [14] explicitly extended this idea to detect concepts encoded in neural networks. The concepts (i.e., perceptible patterns) coded in neural networks could deepen our understanding of DNNs’ operating principles and eventually provide explanations (readable to humans) of DNNs’ decisions. More importantly, it raises the possibility that DNNs could build some abstract codes in hidden layers, some of which could directly be linked to concepts. In this study, we extended this line of studies by seeking abstract codes, which are not necessarily recognizable to humans, in hidden layers functionally linked to DNNs’ decisions. We also explored potential methods of utilizing hidden layers’ abstract codes to make DNNs’ operations/decision-making more reliable and safer.
To make a clear distinction between human recognizable concepts and DNNs’ own abstract codes, we will refer to DNNs’ abstract codes as ‘symbols’ hereafter. Our analysis of hidden layer representations is based on four assumptions. First, symbols ($S_{i}$, abstract codes of DNNs) are encoded across feature maps in each layer. Second, responses of individual hidden layers $H_{i}$ can be described by the sum of symbols $S_{i}$ and residuals $\epsilon$; $H_{i}=S_{i}+\epsilon$. Conversely, we can map individual representations $H_{i}$ into nearest $S_{i}$ by minimizing $\epsilon$. Third, there is a finite number of symbols in DNNs, which is far less than the possible number of hidden layer representations. Fourth, symbols are spatially local and thus need to be extracted from proper receptive fields or regions of interest (ROIs). With these assumptions in mind, we performed unsupervised clustering analysis with hidden layer responses ($H_{i}$) to determine the cluster centers and identify ‘symbols’. Our analysis suggests that these symbols can mediate semantic meanings and allow us to 1) monitor the models’ decision-making process, 2) detect abnormal operations related to adversarial perturbations and out-of-distribution (OOD) examples and 3) temporarily learn about OOD examples. Based on the results, we argue that hidden layers’ responses can be converted to discrete symbols that can help us build more reliable and safer DNNs.
@@ 2 Extracting symbols underlying DNNs’ operations
To test if DNNs learn to use symbols to make decisions, we analyzed ImageNet models’ hidden layer responses to a subset (Mixed_13) of ImageNet [15], previously curated by a python machine learning library [16]. Mixed_13 contains 78 classes out of 1,000 classes of ImageNet, which belong to one of 13 super-classes (Supplementary Table1). Mixed_13 consists of two disjointed ‘training’ and ‘test’ sets. To extract the symbols, we assumed that the symbols associated with DNN’s decision-making would be observed frequently. For instance, if a symbol is associated with a tiger, a picture of a tiger would elicit it frequently. With this assumption in mind, we presented the inputs from Mixed_13 and searched for repeatedly occurring hidden layer responses by using unsupervised clustering analysis.
We note that analyzing hidden layer responses can be challenging for two reasons. First, hidden layer responses are extremely noisy high-dimensional data. A massive number of hidden neurons exist in modern DNNs, and their responses are determined with diverse inputs, making dimension reduction essential (see [17, 18, 13] , for instance). Second, visual scenes often contain multiple objects rather than a single one. For instance, a visual scene may include a tiger or tigers walking along the river, playing with other tigers and/or climbing a tree. Thus, if we want to extract codes related to tigers, it is necessary to isolate regions of interest (ROIs) containing tigers and exclude confounding objects (e.g., tree or river). However, ImageNet training examples are labeled with single class names and do not provide sufficient information to identify ROIs.
@@ 2.1 ROIs Identification
We note that ‘Second Thought Certification’ (STCert) proposed in an earlier study [20] can be used to identify ROIs related to visual objects with high accuracy. STCert is a two-step process (Fig. 1). First, it uses foundational segmentation models [21, 22] to detect ROIs related to DNNs’ predictions. For instance, if a DNN predicts a tiger, STCert asks the segmentation model to return the bounding box (i.e., ROI) of the tiger in the picture. Second, it crops ROIs and uses them to confirm or reject DNNs’ original predictions. STCert provides a more reliable ROI estimates than a single step segmentation, since its segmentation is crosschecked with a classifier. As such, we used STCert to identify ROIs in this study to extract ROIs from randomly chosen 200 training examples for each class in Mixed_13; see Supplementary text A.1 for more detail. Since Mixed_13 contains 78 classes (Supplementary Table1), the total number of training examples is 15,600. Among them, we first selected examples, on which DNNs’ predictions were correct to minimize the bias induced by the incorrect predictions/decisions of DNNs (i.e., incorrect mapping between ROI and the inputs’ labels). For some of the images, STCert did not find ROIs or found multiple ROIs, which means the actual number of ROIs are determined by a few factors, not just the number of training examples.
@@ 2.2 Generating hidden layer activity vectors
Once ROIs were identified, we sampled activations ($H_{i}$) of hidden neurons, whose receptive fields correspond to ROIs. To reduce noise and filter out non-essential information, we used ROI-pooling [19] to convert hidden layer activations into 3-by-3 activation maps. That is, the size of hidden layer activation maps can be arbitrary depending on the size of objects, but ROI-pooling always turns them into 3-by-3 activation maps. It should be noted that individual activations in 3-by-3 activation maps reflect pixels at different spatial locations. As we assumed that symbols are collectively encoded neurons that share the same receptive field, we built vector codes (referred to as activity vectors below) by collecting activations at the same location in the maps across channels. For instance, there are 512 channels in ResNet18’s layer 4, and we have 9 location-specific activation maps in each channel. In this case, we aggregated activations from the same location in the activation map from all 512 channels and then created 9 activity vectors, each of which contains 512 components, from the layer per image.
In this study, we analyzed 4 hidden layers of popular DNNs, ResNet18, ResNet50, VGG19 and DenseNet121 [23, 24, 25]. ResNet and DenseNet consist of 4 composite blocks, and their outputs were analyzed. VGG19 contains 5 max pooling layers, whose outputs roughly correspond to ResNet’s composite block’s outputs. We selected the last 4 max-pooling layers and analyzed them, as they match the layers of ResNet and DenseNet. We used Pytorch [26] and Timm [27] python deep learning libraries to implement these models. A workstation with Intel Core9 CPU and RTX 4090 was used for all our experiments and analyses.
@@ 2.3 Clustering analysis of hidden layer activity vectors
We turned to unsupervised clustering analysis to identify representative activity vectors equivalent to symbols in this study. We also note that many activity vectors contain low values. With these low-value activity vectors, clustering algorithms consider the zero-valued vector as the most prominent cluster center, which is not informative. Thus, we removed the vectors consisting of values lower than the mean activity value estimated on each layer. Next, we used UMAP analysis [28, 29] to further reduce them to 3-dimensional vectors; see Table 1 for the parameters used in the analysis. These three dimensional vectors were analyzed to detect symbols via clustering. It should be noted that the exact number of clusters existing in the activity vector space are unknown, and the result of clustering analysis strongly depends on the number of predefined clusters. To address this issue, we used X-means clustering [30] to automatically discover an optimal number of clusters (i.e., the number of symbols). We note that X-means is an extension of Kmeans clustering, which uses Bayesian Information Criterion (BIC) to determine the optimal number of cluster, and a python package ‘pyclustering’ [31] was used to implement X-means in this study. In our analysis, we restrict the maximum number of classes to 1,000. Table 2 shows the identified number of clusters in 4 layers of DNNs tested. As shown in the table, the optimal numbers of clusters in early layers are smaller than 1,000, but those in the late layers hit the maximum value, suggesting that homogeneity of activity vectors increased, as layers go deeper.
@@ 3.1 Links between symbols and semantic meanings of inputs
$CM(i,j)=\begin{cases}CM(i,j)+1&\text{if $i=S_{mn}$, $j=l_{m}$},\text{where }n=1,...,9\\ CM(i,j)&\text{otherwise}\par\end{cases}$ (1)
, where $CM(i,j)$ is the component of Correlation map at $i^{th}$ row and $j^{th}$ column; $S_{mn}$ denotes the $n^{th}$ (out of 9) symbol related to $m^{th}$ ROI; where $l_{m}$ is the label of $m^{th}$ ROI (i.e., $img_{k}$).
If a symbol is associated with a specific class, it would appear frequently when the class object is presented, but it would not appear when other class objects are presented. Fig. 2 shows the correlation $CM(i,j)$ between 100 randomly chosen symbols from 4 layers of ResNet18 and 78 classes. $Y$-axis denotes 100 randomly chosen symbols, and $x$-axis denotes the label (i.e., class). We found two types of symbols. First, some symbols were associated with a broad range of examples. Second, some symbols were selectively linked to specific classes. Symbols specific to classes were rare in early layers but became more common in late layers. We found equivalent results in ResNet50, VGG19 and DenseNet121 (Supplementary Fig. 1). In addition, we note that many symbols can be merged together (Supplementary Fig. 2) and thus, did not increase the maximum number of clusters above 1000. These results suggest that some symbols could be directly linked to inputs’ labels, and thus, one could infer the labels from their associated symbols.
@@
To test this possibility, we obtained the symbols associated with test examples of Mixed_13 and asked if they can be used to predict the labels of inputs (i.e., ROIs). In this experiment, we collected ROIs from all test images whether or not DNNs make correct predictions. From these ROIs, we collected activity vectors and obtained the test symbols (i.e., the nearest cluster centers). These ‘test symbols’ were used to predict the likelihood of each class by using $CM(i,j)$ (Eq. 1). As we aimed to compute the probability of the class, the correlation map $CM(i,j)$ was normalized along with a row (i.e., over classes) using SoftMax (2). By using the normalized Correlation Map $P(i,j)$, we calculated the probability $P(i,j)^{m}$ that $m^{th}$ ROI belonged to the class $j$ when the symbol $S_{i}$ was observed (Eq. 2). That is, symbol indices were used as hash keys for a look-up table.
$P(i=S_{i},j)^{m}=\frac{e^{C(i=S_{i},j)}}{\sum_{n=1}^{78}e^{C(i=S_{i},n)}}$ (2)
As we obtained 9 symbols from each ROI, we took average probability over 9 (Eq. 3) and then determine the most probable class as the prediction $Pr_{m}$ (Eq. 4).
@@
$Pr_{m}=\operatorname*{argmax}_{j}P(j)^{m}$ (4)
Fig. 3 shows the accuracy of the symbol-based predictions depending on the layer. We note that the prediction accuracy is around 80% in layer 4 and much lower in earlier layers. These accuracies cannot be directly compared to the models’ accuracy because the choice is limited to one out of 78 classes, not 1000 classes. However, the accuracy in layer 4 and earlier layers were significantly higher than chance (1/78), supporting that the identified symbols were associated with semantic meanings.
If symbols are associated with semantic meanings, they could also be associated with DNNs’ predictions. We tested this possibility by correlating the symbols and DNNs’ predictions on Mixed_13 training examples (once again, 200 examples per class) instead of inputs’ labels. As stated above, we used the symbols to infer their predictions on test examples and found that the symbols are predictive of DNNs’ answers on test examples (Fig. 3B), raising the possibility that DNNs could utilize symbols to make decisions. The observed link between symbols and DNNs’ decisions leads us to hypothesize that symbols can be used to augment DNNs’ decision-making process. Below we propose potential ways to make DNNs’ operations more reliable using symbols.
Figure 3: Symbol-based predictions. (A), Symbol-based prediction on the labels of test examples of Mixed_13. (B), Symbol-based prediction on the DNNs’ answers.

### v2

@@ Abstract
Deep learning (DL) enables deep neural networks (DNNs) to automatically learn complex tasks or rules from given examples without instructions or guiding principles. As we do not engineer DNNs’ functions, it is extremely difficult to diagnose their decisions, and multiple lines of studies proposed to explain the principles of their operations. Notably, one line of studies suggests that DNNs may learn concepts, the high level features that are recognizable to humans. In this study, we extend this line of studies and hypothesize that DNNs can develop abstract codes that can be used to augment DNNs’ decision-making. To address this hypothesis, we combine foundation segmentation models and unsupervised learning to extract internal codes and identify potential use of abstract codes to make DL’s decision-making more reliable and safer.
@@ 1 Introduction
Deep learning (DL) can automatically wire deep neural networks (DNNs) to perform complex tasks, and DNNs’ learning capability allows them to outperform humans in some tasks [1, 2, 3]. Notably, instead of expert instructions or feature engineering, DL requires only training examples to solve highly complex problems. However, as DNNs’ structures are automatically determined without guiding principles, we do not fully understand their decision-making process. That is, DNNs may be useful black box decision-makers, but without safety measures, deploying them in high-stakes domains (e.g., autonomous cars) comes with unpredictable risks [4, 5]. For the moment, the current generation of DL does not warrant essential safety to be deployed in safety-critical or high-stakes domains. To address this shortcoming, earlier studies sought potential methods to decode and explain DNNs’ decisions. There are multiple lines of research in explainable AI (e.g., see Refs. [6, 7, 8, 9, 10]), but in our study we pay special attention to the one probing correlations between hidden layer representations and inputs or between hidden layer representations and outputs (i.e., decisions). Feature visualization [11] and network dissection [12] sought direct links between inputs and hidden layers, while linear probes investigated correlations between outputs and hidden layers [13, 14].
@@
It is notable that the mapping between hidden layers and visual features in input layers, in principle, establishes links between high level patterns recognizable to humans (often referred to as ‘concepts’) and hidden neurons. The Test with Concept Activation Vectors [14] explicitly extended this idea to detect concepts encoded in neural networks. The concepts (i.e., perceptible patterns) encoded in neural networks could deepen our understanding of DNNs’ operating principles and eventually provide explanations (readable to humans) of their decisions. More importantly, it raises the possibility that DNNs could build some abstract codes in hidden layers, some of which could directly be linked to concepts. In this study, we extend this line of studies by seeking abstract codes, which are not necessarily recognizable to humans, in hidden layers functionally linked to DNNs’ decisions. To make a clear distinction between human recognizable concepts (inputs’ features) and DNNs’ own abstract codes, we will refer to DNNs’ abstract codes as ‘symbols’ hereafter.
Our analysis is inspired by two observations in the literature. First, the hidden layers encode distinct contextual information. Earlier layers encode low-level features such as texture and shapes, whereas late layers encode high-level semantic features. Second, the penultimate layer’s outputs are clustered together according to the labels of inputs. As all hidden layers are trained simultaneously via the same algorithm, it seems natural to assume that all hidden layers may share the same operating principles. More specifically, we assume that hidden layers map inputs with the same contextual information into neighboring outputs, as the penultimate layer maps inputs belonging to the same class onto the same clusters. The penultimate layer encodes the labels, and the earlier hidden layers encode different contexts currently unknown to us. If these functional clusters reflecting contextual information exist in the hidden layers, identifying them can conversely reveal how inputs are progressively encoded. That is, we could uncover the symbols using these functional clusters.
With this possibility in mind, we seek the functional clusters in the hidden layers to obtain the symbols, the abstract codes of DNNs. We note that visual scenes generally consist of numerous distinct objects, and thus, it is necessary to separate them. To this end, we use segmentation models to locate visual objects of interest and analyze the responses of hidden neurons whose receptive fields cover the locations. Our analysis suggests that these symbols can mediate semantic meanings and allow us to 1) monitor the models’ decision-making process, 2) detect abnormal operations related to adversarial perturbations and out-of-distribution (OOD) examples and 3) temporarily learn about OOD examples. Based on these results, we propose that hidden layers’ responses can be converted to discrete symbols that can help us build more reliable and safer DNNs.
@@ 2 Extracting symbols underlying DNNs’ operations
To test whether DNNs learn to use symbols to make decisions, we analyzed 5 ImageNet models’ hidden layers using a subset (Mixed_13) of ImageNet [15], previously curated by a python machine learning library [16], and the hidden layer responses of ResNet50 [17] trained on the Oxford-IIIT Pet dataset [18] . Mixed_13 contains 78 classes out of 1,000 classes of ImageNet, which belong to one of 13 super-classes (Supplementary Table 1). Mixed_13 consists of two disjointed ‘training’ and ‘test’ sets. Oxford-IIIT Pet dataset contains 37 classes of cat and dog breeds. To extract the symbols, we assumed that the symbols associated with DNN’s decision-making would be observed frequently. For instance, if a symbol is associated with a tiger, a picture of a tiger would elicit the symbol frequently. With this assumption in mind, we presented the inputs from Mixed_13 and searched for repeatedly occurring hidden layer responses by using unsupervised clustering analysis.
It should be noted that analyzing hidden layer responses can be challenging for two reasons. First, hidden layer responses are extremely noisy high-dimensional data. A massive number of hidden neurons exist in modern DNNs, and their responses are determined by diverse inputs, making dimension reduction essential (see [19, 20, 13] , for instance). Second, visual scenes often contain multiple objects rather than a single one. For instance, a visual scene may include a tiger or tigers walking along the river, playing with other tigers and/or climbing a tree. Thus, to extract codes related to tigers, we need to isolate regions of interest (ROIs) containing tigers and exclude confounding objects (e.g., tree or river). However, ImageNet training examples are labeled with single class names and do not provide sufficient information to identify ROIs.
@@ 2.1 ROIs Identification
We note that ‘Second Thought Certification’ (STCert) proposed in an earlier study [22] can be used to identify ROIs related to visual objects with high accuracy. STCert is a two-step process (Fig. 1). First, it uses foundational segmentation models [23, 24] to detect ROIs related to DNNs’ predictions. For instance, if a DNN predicts a tiger, STCert asks the segmentation model to return the bounding box (i.e., ROI) of the tiger in the picture. Second, it crops ROIs and uses them to confirm or reject DNNs’ original predictions. STCert provides a more reliable ROI estimates than a single step segmentation, since its segmentation is crosschecked with a classifier. That is, STCert relies on ‘self-consistency’. As such, we used STCert to identify ROIs in this study to extract ROIs from randomly chosen 200 training examples for each class in Mixed_13; see Supplementary text for more detail. Since Mixed_13 contains 78 classes (Supplementary Table 1), the total number of training examples is 15,600. Among them, we first selected examples, on which DNNs’ predictions were correct to minimize the bias induced by the incorrect predictions/decisions of DNNs (i.e., incorrect mapping between ROI and the inputs’ labels). For some of the images, STCert did not find ROIs or found multiple ROIs, which means the actual number of ROIs is determined by a few factors, not just the number of training examples.
@@ 2.2 Generating hidden layer activity vectors
Once ROIs were identified, we sampled activations ($H_{i}$) of hidden neurons, whose receptive fields correspond to ROIs. To reduce noise and filter out non-essential information, we used ROI-pooling [21] to convert hidden layer activations into 3-by-3 activation maps. That is, the size of hidden layer activation maps can be arbitrary depending on the size of objects, but ROI-pooling always turns them into 3-by-3 activation maps. It should be noted that individual activations in 3-by-3 activation maps reflect pixels at different spatial locations. As we assumed that symbols are collectively encoded neurons that share the same receptive field, we built vector codes (referred to as activity vectors below) by collecting activations at the same location in the maps across channels. For instance, there are 512 channels in ResNet18’s layer 4, and we have 9 location-specific activation maps in each channel. Here, we aggregated activations from the same location in the activation map from all 512 channels and then created 9 activity vectors, each of which contains 512 components from the layer per image.
In this study, we analyzed 4 hidden layers of popular DNNs, ResNet18, ResNet50, VGG19, DenseNet121, and Vision Transformers (ViT) [17, 25, 26, 27]. ResNet and DenseNet consist of 4 composite blocks, and their outputs were analyzed. VGG19 contains 5 max pooling layers, whose outputs roughly correspond to ResNet’s composite block’s outputs. We selected the last 4 max-pooling layers and analyzed them, as they match the layers of ResNet and DenseNet. For ViT consisting of 12 layers, we chose the 4 layers (layer ids: 3, 6, 9, 12) . We used Pytorch [28] and Timm [29] python deep learning libraries to implement these models. A workstation with Intel Core9 CPU and RTX 4090 was used for all our experiments and analyses.
@@ 2.3 Clustering analysis of hidden layer activity vectors
We turned to unsupervised clustering analysis to identify representative activity vectors equivalent to symbols in this study. We also note that many activity vectors contain low values. With these low-value activity vectors, clustering algorithms consider the zero-valued vector as the most prominent cluster center, which is not informative. Thus, we removed the vectors consisting of values lower than the mean activity value estimated on each layer. Next, we used UMAP analysis [30, 31] to further reduce them to 3-dimensional vectors; see Table 1 for the parameters used in the analysis. These three dimensional vectors were analyzed to detect symbols via clustering. It should be noted that the exact number of clusters existing in the activity vector space are unknown, and the result of clustering analysis strongly depends on the number of predefined clusters. To address this issue, we used X-means clustering [32] to automatically discover an optimal number of clusters (i.e., the number of symbols). We note that X-means is an extension of Kmeans clustering, which uses Bayesian Information Criterion (BIC) to determine the optimal number of clusters, and a python package ‘pyclustering’ [33] was used to implement X-means in this study. In our analysis, we restrict the maximum number of classes to 1,000. Table 2 shows the identified number of clusters in 4 layers of DNNs tested. As shown in the table, the optimal numbers of clusters in the early layers are smaller than 1,000, but those in the late layers hit the maximum value, suggesting that homogeneity of activity vectors decreases, as the layers go deeper.
@@ 3.1 Links between symbols and semantic meanings of inputs
$CM(i,j)=\begin{cases}CM(i,j)+1&\text{if $i=S_{mn}$, $j=l_{m}$}\\ CM(i,j)&\text{otherwise}\par\end{cases}$ (1)
, where $n=1,...,9$, and $CM(i,j)$ is the component of Correlation map at $i^{th}$ row and $j^{th}$ column; $S_{mn}$ denotes the $n^{th}$ (out of 9) symbol related to $m^{th}$ ROI; where $l_{m}$ is the label of $m^{th}$ ROI (i.e., $img_{k}$).
If a symbol is associated with a specific class, it would appear frequently when the class object is presented, but it would not appear when other class objects are presented. Fig. 2 shows the correlation $CM(i,j)$ between 100 randomly chosen symbols from 4 layers of ResNet18 and 78 classes. $Y$-axis denotes 100 randomly chosen symbols, and $x$-axis denotes the label (i.e., class). We found two types of symbols. First, some symbols were associated with a broad range of examples. Second, some symbols were selectively linked to specific classes. Symbols specific to classes were rare in early layers but became more common in late layers. We found equivalent results in ResNet50, VGG19 and DenseNet121 (Fig. 3). Interestingly, we note that many symbols can be merged together (Fig. 4), suggesting that many symbols can be redundant. Thus, we did not increase the maximum number of clusters above 1000. These results suggest that some symbols could be directly linked to the labels of inputs, and thus, one could infer the labels from their symbols in the hidden layers.
@@
Figure 3: The correlations between symbols and classes. The top, middle and bottom rows show the correlations observed from ResNet50, VGG19 and DenseNet121. $y$-axis denotes the indices of symbols, $x$-axis denotes the class. (A)-(D), the correlations observed in layers 1-4, respectively.
Figure 4: The correlations between symbols and classes in ResNet18. These plots show the correlation maps between 200 consecutive symbols and the labels of inputs observed in layer 1 (A), layer 2 (B), layer 3 (C) and layer 4 (D). We note that some consecutive symbols are correlated with the same classes and thus can be merged together, especially in layer 4.
To test this possibility, we obtained symbols associated with test examples of Mixed_13 and asked if they could be used to predict the labels of inputs (i.e., ROIs). In this experiment, we collected ROIs from all test images whether or not DNNs make correct predictions. From these ROIs, we collected activity vectors and obtained the test symbols by identifying the nearest cluster centers. These ‘test symbols’ were used to predict the likelihood of each class by using $CM(i,j)$ (Eq. 1). As we aimed to compute the probability of the class, the correlation map $CM(i,j)$ was normalized along with a row (i.e., over classes) using SoftMax (2). By using the normalized Correlation Map $P(i,j)$, we calculated the probability $P(i,j)^{m}$ that $m^{th}$ ROI belonged to the class $j$ when the symbol $S_{i}$ was observed (Eq. 2). That is, symbol indices were used as hash keys for a look-up table.
$P(i=S_{i},j)^{m}=\frac{e^{CM(i=S_{i},j)}}{\sum_{n=1}^{78}e^{CM(i=S_{i},n)}}$ (2)
As we obtained 9 symbols from each ROI, we took average probability over 9 (Eq. 3) and then determined the most probable class as the prediction $Pr_{m}$ (Eq. 4).
@@
$Pr_{m}=argmax_{j}P(j)^{m}$ (4)
Fig. 5 shows the accuracy of the symbol-based predictions for all 5 ImageNet models and ResNet50 trained Oxford-IIT PET dataset. We note that the prediction accuracy is around 80% in layer 4 (i.e., $5^{th}$ Max-pooling in VGG19 and $12^{th}$ embedding layer in ViT) and much lower in the earlier layers. These accuracies of ImageNet model cannot be directly compared to the models’ accuracy because the choice is limited to one out of 78 classes, not 1000 classes. However, the accuracy in layer 4 and the earlier layers were significantly higher than chance (1/78), supporting that the identified symbols were associated with semantic meanings. ResNet50 trained on Oxford-IIT PET dataset also show a similar level of accuracy.
If symbols are associated with semantic meanings, they could also be associated with DNNs’ predictions. We tested this possibility by correlating the symbols and DNNs’ predictions on Mixed_13 training examples (once again, 200 examples per class) instead of inputs’ labels. As stated above, we used the symbols to infer their predictions on test examples and found that the symbols are predictive of DNNs’ answers on test examples (Fig. 5B). We made equivalent observation from ResNet50 trained on Oxford-IIT PET dataset. These observations raise the possibility that DNNs could utilize symbols to make decisions.
Figure 5: Symbol-based predictions. (A), Symbol-based prediction on the labels of test examples of Mixed_13. (B), Symbol-based prediction on the DNNs’ answers. PET denotes ResNet50 trained on Oxford-IIT PET. L1, L2, L3 and L4 denote the analyzed layers. For ViT, they denote $3^{rd},6^{th},9^{th}$, and $12^{th}$ embedding layer. For VGG19, they denote $2^{nd},3^{rd},4^{th}$ and $5^{th}$ Max-pooling layers. For ResNet and DenseNet, they denote all 4 layer blocks.
We also tested the predictive power of symbols depending on the number of symbols. Specifically, we set a number of symbols from the 4 layers of ResNet18 to a specific number by using KMeans clustering instead of X-Means clustering. Then, we measured the predictive power of extracted symbols on labels of test examples by evaluating the accuracy of symbol-based predictions. Specifically, we tested 5 different numbers of symbols, which are 500, 1000, 1500, 2000 and 2500. As shown in Table 3, the predictive power (i.e., accuracy) of symbols does not monotonically increase, as the number of symbols increases, suggesting that the predictive power of symbols is not sensitive to the exact number of symbols.
The observed link between symbols and DNNs’ decisions leads us to hypothesize that symbols can be used to augment DNNs’ decision-making process. Below we propose potential ways to make DNNs’ operations more reliable using symbols.
Table 3: Symbol-based predictions depending on the number of symbols. Instead of $x$-means, we set the number of symbols to be 500, 1000, 1500, 2000 and 2500. With the number of symbols fixed, we estimated the accuracy of the symbol-based predictions in 4 layers.
Layer 500 1000 1500 2000 2500
1 0.218 0.219 0.23 0.229 0.228
2 0.264 0.27 0.285 0.28 0.283
3 0.431 0.472 0.478 0.486 0.48
4 0.782 0.805 0.812 0.809 0.819